### v1

# Proper maps of annuli

Abstract.
We study rational proper holomorphic maps of annuli in complex euclidean spaces, that is, domains with $U(n)$ as the automorphism group. By the Hartogs phenomenon and a result of Forstnerič, such maps are always rational and extend to proper maps of balls. We first prove that a proper map of annuli from $n$ dimensions to $N$ dimensions where $N<\binom{n+1}{2}$ is always an affine embedding. This inequality is sharp as the homogeneous map of degree 2 satisfies $N=\binom{n+1}{2}$. Next we find a necessary and sufficient condition for a map to be homogeneous: A proper map of annuli is homogeneous if and only if its general hyperplane rank, the affine dimension of the image of a general hyperplane, is exactly N1N-1. This result is really a result classifying homogeneous proper maps of balls. A homogeneous proper ball map takes spheres centered at the origin to spheres centered at the origin. We show that if a proper ball map has general hyperplane rank N1N-1 and takes one sphere centered at the origin to a sphere centered at the origin, then it is homogeneous. Another corollary of this result is a complete classification of proper maps of annuli from dimension 2 to dimension 3.
Key words and phrases:
rational sphere maps, proper holomorphic mappings, maps of annuli, Segre varieties, affine dimension
2020 Mathematics Subject Classification:
32H35, 32A08, 32H02
The second author was in part supported by Simons Foundation collaboration grant 710294.
1. Introduction
Let $\mathbb{B}_{n}\subset\mathbb{C}^{n}$ denote the unit ball. A well-studied problem in complex analysis is to attempt to classify the proper holomorphic maps $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. If $n=N=1$, then by a classical theorem of Fatou [fatou-1923-fonctions], $f$ is a finite Blaschke product and hence rational. If $n=N>1$, then by a theorem of Alexander [alexander-1977-proper], $f$ is an automorphism of the unit ball. No maps exist if $N<n$ for elementary reasons. Dor and Løw [dor-1990-proper, low-1985-embeddings] show that if $N>n>1$, there exist non-rational maps, even in the case of codimension 1, that is, $N=n+1$. If in the case $n>1$ one assumes some regularity on the boundary, then by a theorem of Forstnerič [forstneric-1989-extending], all such $f$ are rational with the degree bounded by an integer depending on $n$ and $N$. Faran [faran-1982-maps] showed that proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ that are $C^{3}$ up to the boundary are equivalent to one of four polynomial maps. For maps that are $C^{3}$ up to the boundary, Webster [webster-1979-mapping] showed that for $n\geq 3$, there is no map from $\mathbb{B}_{n}$ to $\mathbb{B}_{n+1}$ that is not equivalent up to automorphisms (spherically equivalent) to the linear embedding $z\oplus 0$. Faran [faran-1986-linearity] proved the same conclusion (only the linear embedding) for maps real-analytic up to the boundary for all $N$ up to N=2n2N=2n-2. Huang [huang-1999-linearity] improved the regularity to $C^{2}$. For N=2n1N=2n-1, the Whitney map $(z^{\prime},z_{n})\mapsto(z^{\prime},z_{n}\otimes z)$ is not spherically equivalent to the linear embedding. The map $z^{\otimes d}$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$. This map can be symmetrized (as $z^{\otimes d}$ contains some duplicate entries) via a unitary mapping to what we will call $H_{d}$ direct sum a number of zero components. Rudin [rudin-1984-homogeneous] proved that every homogeneous proper ball map is, up to a unitary, $H_{d}\oplus 0$. See D’Angelo [dangelo-1988-polynomial] for a simpler proof of this fact. In general, see the books by D’Angelo [dangelo-1993-several, dangelo-2019-hermitian, dangelo-2021-rational] for a more thorough treatment of proper maps of balls.
The ball can be regarded as the strongly pseudoconvex domain with the most symmetry. Let us consider then a non-pseudoconvex domain with a lot of symmetry, the annulus. Write $\mathbb{A}_{n,r}=\mathbb{B}_{n}\setminus\overline{r\mathbb{B}_{n}}=\{z\in\mathbb{C}^{n}:r<\mathopen{}\mathclose{{\left\lVert z}}\right\rVert<1\}$ for an annulus, where we will always assume that $0<r<1$. The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group as long as $n>1$. When $n=1$ there are also the automorphisms $e^{i\theta}\frac{r}{z}$. The same classical argument to classify the automorphism group $\mathbb{A}_{1,r}$ can be extended to show that in one dimension a proper holomorphic map $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ exists only if $R=r^{d}$ for some $d\in\mathbb{N}$ in which case it is either equal to $e^{i\theta}z^{d}$ or $r^{d}e^{i\theta}z^{-d}$. In dimension $n\geq 2$, due to the inner pseudoconcave boundary, via the Hartogs phenomenon and the theorem of Forstnerič on rationality of proper ball maps, all proper holomorphic maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are rational, see Theorem˜2.2, but they can be more complicated. In fact, holomorphic proper maps of annuli are in one to one correspondence with rational proper maps of balls that take a sphere centered at zero to another sphere centered at zero. See [helal-2025-proper] for more details.
For $z\in\mathbb{C}^{n}$, if $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is constant, then
(1) $\mathopen{}\mathclose{{\left\lVert H_{d}(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z^{\otimes d}}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2d}$
is constant. Hence, $H_{d}$ is a map that takes any sphere of radius $r$ centered at the origin to a sphere of radius $r^{d}$ centered at the origin. Moreover, it is not difficult to show that it takes $\mathbb{A}_{n,r}$ (for any $0<r<1$) properly to $\mathbb{A}_{N,R}$, where $N$ is the number of components of $H_{d}$ and $R=r^{d}$. Similarly, if we have any two proper maps $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{1},R_{1}}$ and $G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{2},R_{2}}$ and $t\in(0,1)$, then $\sqrt{1-t}\,F\oplus\sqrt{t}\,G\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N=N_{1}+N_{2}$ and $R^{2}=(1-t)R_{1}^{2}+tR_{2}^{2}$. We call such a weighted direct sum a juxtaposition. So juxtapositions of proper maps of annuli are proper maps of annuli as long as the source annulus is the same. Based on the one-dimensional result, noting that $z^{\otimes d}$ is the correct multidimensional analogue of $z^{d}$, one could (incorrectly) conjecture that all such proper maps are juxtapositions of homogeneous maps. However, other maps do exist, see Example˜2 and more generally [helal-2025-proper].
Our goal is to start a classification of proper maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$. We say $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ and $g\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ are unitarily equivalent (or simply equivalent) if there exist unitary matrices $V\in U(n)$ and $W\in U(N)$ so that
(2) $f=WgV.$
As we noted, $U(n)$ is the holomorphic automorphism group of $\mathbb{A}_{n,r}$. If $n=N>1$, then a proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ exists only if $r=R$ and $f$ is unitarily equivalent to the identity $z$, see Corollary˜2.6. If $N>n$, then an affine embedding exists whenever $r\leq R<1$, but is not necessarily $z\oplus 0$ due to the different automorphism group. In particular, $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1, then $f$ is unitarily equivalent to the affine embedding
(3) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0,$
where the last 0 has Nn1N-n-1 terms. That is, we shift the linear embedding by adding the proper constant in the $(n+1)$th component. See Proposition˜2.7.
As we noted above, for sufficiently smooth proper maps between balls, $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ no maps other than the linear embedding exist for N<2n1N<2n-1, and then when N=2n1N=2n-1, the Whitney mapping exists and is not spherically equivalent to a linear embedding. Curiously, for annulus maps, this gap is much longer and goes until $N=\binom{n+1}{2}$.
Theorem 1.1.
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map and $2\leq n<N<\binom{n+1}{2}$. Then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$. The bound on $N$ is sharp as the homogeneous map $H_{2}$ takes the annulus $\mathbb{A}_{n,r}$ properly to $\mathbb{A}_{\binom{n+1}{2},r^{2}}$.
As homogeneous maps play an outsize role in the study of proper maps of annuli we next find a property that determines a map to be homogeneous. The framework that we find applies more broadly to proper maps of balls and so we state it first in that setting. First we define a certain invariant for proper mappings that has, in various guises, appeared in the study of ball maps. First let $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$ be any holomorphic map. If $X\subset\mathbb{C}^{N}$, write $\operatorname{aff}(X)$ for the affine hull of $X$, that is, the smallest affine subspace of $\mathbb{C}^{N}$ containing $X$. Define
(4) $k_{f}\overset{\text{def}}{=}\max\mathopen{}\mathclose{{\left\{\dim\operatorname{aff}\bigl(f(H\cap U)\bigr):H\subset\mathbb{C}^{n}\text{ is an affine hyperplane}}}\right\}.$
We call $k_{f}$ the general hyperplane rank. One could think of $k_{f}$ as a certain measure of complexity. It is not difficult to see that $k_{f}$ is attained on a generic hyperplane, and it is in general quite easy to compute, especially for rational maps. While for an arbitrary holomorphic map one generally expects $k_{f}=N$, via a Segre-variety argument, a rational proper map of balls has $k_{f}\leq N-1$. The idea that Dor [dor-1990-proper] used to show a proper map is not rational is to, in other words, show that the constructed map has $k_{f}=N$ as any rational ball map has $k_{f}=N-1$. If $k_{f}=N-1$, then we obtain a map of Grassmannians. That technique was used for example by Faran to prove his classification of maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The number $k_{f}$ is related to the $X$-variety that Forstnerič used to show the rationality result. In particular, for a rational map, if the dimension of the generic fiber of the $X$-variety is $\ell$, then $k_{f}=N-\ell-1$. D’Angelo [dangelo-2003-Xvariety] has studied the $X$-variety for rational maps and developed a technique of quickly computing fibers of the $X$-variety as the nullity of a certain matrix $C(w)$. The nullity of $C(w)$ for a generic $w$ is then the $\ell$. A related concept is the embedding dimension $N_{f}$, that is, the dimension of the smallest affine subspace $S$ containing $f(\mathbb{B}_{n})$ (or $f(\mathbb{A}_{n,r})$). In other words, $N_{f}=\dim\operatorname{aff}f(\mathbb{B}_{n})$. In either case, restricting to $S$ on the target we obtain a proper map between balls or annuli with target dimension $N_{f}$. While for rational proper maps, often $k_{f}=N_{f}-1$, it is not always true. Example˜2 mentioned above has $k_{f}=N_{f}-3$. A juxtaposition of two rational maps also has $k_{f}\leq N_{f}-2$, see Proposition˜2.12. What we prove is that for proper maps of annuli, $k_{f}=N_{f}-1$ is precisely the condition that characterizes homogeneous maps $H_{d}$.
Theorem 1.2.
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map. Then $k_{f}=N-1$ if and only if $N=\binom{n+d-1}{d}$, $R=r^{d}$, and
(5) $f=UH_{d},$
for some $d\in\mathbb{N}$ and a unitary $U\in U(N)$.
One direction of this theorem is obvious, it is that $k_{f}=N-1$ implies that $f$ is homogeneous that is remarkable. The proof relies on a result from [helal-2025-proper], where the authors showed that a rational map that takes infinitely many spheres centered at the origin to spheres centered at the origin is a juxtaposition of homogeneous maps. We note that if we wish to state the theorem in terms of $N_{f}$, then we ought to say that $k_{f}=N_{f}-1$ if and only if $f=U(H_{d}\oplus 0)$. We note that $k_{f}=N-1$ implies that $N=N_{f}$ for any proper rational map of annuli (or balls).
Let us state a version of the forward direction of this result for proper maps of balls. Clearly $H_{d}$ is a proper ball map that takes infinitely many spheres centered at the origin to spheres centered at the origin. With the assumption that $k_{f}=N-1$, to show that $f$ is homogeneous, it is sufficient to check that a map takes just one inner sphere centered at the origin to a sphere centered at the origin.
Theorem 1.3.
Suppose $n\geq 2$ and $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ is a rational proper map such that $k_{f}=N-1$ and $f(rS^{2n-1})\subset RS^{2N-1}$ for some $r,R\in(0,1)$. Then there exists a unitary map $U$ and an integer $d$ such that $f=UH_{d}$.
The proof follows by using the $k_{f}=N-1$ condition to show that $f$ must in fact take infinitely many spheres centered at zero to spheres centered at zero.
As a corollary to Theorem˜1.2 and to the result of Faran classifying proper maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$, we completely classify proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$.
Theorem 1.4.
Suppose $f\colon\mathbb{A}_{2,r}\to\mathbb{A}_{3,R}$ is a proper holomorphic map. Then $f$ is unitarily equivalent to exactly one of the following two maps:
  1. (i)
    The affine embedding
    (6) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{1},\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{2},\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}}}\right),$
    where $R\geq r$. Note that $N_{f}=2$ and $k_{f}=1$.
  2. (ii)
    The homogeneous map $H_{2}$
    (7) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(z_{1}^{2},\sqrt{2}\,z_{1}z_{2},z_{2}^{2}}}\right),$
    where $R=r^{2}$. Note that $N_{f}=3$ and $k_{f}=2$.
Let us give the organization of this paper. In Section˜2, we state and prove the basic results for maps of annuli. In Section˜3, we prove the first gap theorem, that is, Theorem˜1.1. In Section˜4, we prove the classification of homogeneous maps in terms of $k_{f}$, that is, Theorem˜1.2. We also prove the classification of annuli maps in dimension 2 and 3, that is, Theorem˜1.4. Finally, in Section˜5, we give some further remarks and open questions.
2. Basic results on proper maps of annuli
Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper mappings of annuli in one dimension.
Proposition 2.1.
Suppose $f\colon\mathbb{A}_{1,r}\to\mathbb{A}_{1,R}$ is a proper holomorphic map. Then there is some $d\in\mathbb{N}$ and $\theta\in\mathbb{R}$ so that $R=r^{d}$ and $f(z)=e^{i\theta}z^{d}$ or $f(z)=r^{d}e^{i\theta}z^{-d}$.
Sketch of proof.
Let us sketch the proof as it may not be easy to find, but it is almost identical to classifying biholomorphisms of annuli. First using an inversion and scaling and the maximum principle, we may assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner. The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and goes to 0 on the outer circle and 1 on the inner circle. Hence it is equal to $\frac{\log\mathopen{}\mathclose{{\left\lvert z}}\right\rvert}{\log r}$, and we get $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{\frac{\log R}{\log r}}$. So $f(z)=e^{i\theta}z^{d}$ and $R=r^{d}$ for some $d\in\mathbb{N}$. ∎
In [helal-2025-proper], the authors have proved, using the theorem of Forstnerič [forstneric-1989-extending] as a key ingredient, that all annulus maps are rational when $n\geq 2$. For completeness we reproduce it here.
Theorem 2.2 ([helal-2025-proper]).
Suppose $n\geq 2$ and $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map. Then $f$ is a rational map of degree bounded by some expression $D(n,N)$. Moreover, $f$ extends to a holomorphic map on $\mathbb{B}_{n}$ and gives a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ and also a proper map of $r\mathbb{B}_{n}$ to $R\mathbb{B}_{N}$.
Unlike in the ball case, no hypothesis on the boundary regularity is needed due to the pseudoconcavity of the inner boundary. The best expression for $D$ is not known, though D’Angelo has conjectured [dangelo-2003-sharp] that it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 is $n=2$.
Sketch of proof.
As $n\geq 2$, the map $f$ extends holomorphically through the inner ball via Hartogs phenomenon and the image is contained in $\mathbb{B}_{N}$ by the maximum principle. It must take the inner sphere to the inner sphere as it takes the pseudoconcave part of the boundary of the source to the pseudoconcave part of the boundary of the target. By the theorem of Forstnerič, $f$ restricted to the inner ball $r\mathbb{B}_{n}$ is a proper rational map to $R\mathbb{B}_{N}$ and gives the degree bound. That $f$ is a proper map of $\mathbb{B}_{n}$ to $\mathbb{B}_{N}$ follows since we must take the pseudoconvex part of the boundary to the pseudoconvex part. ∎
Via a theorem of Cima–Suffridge [cima-1990-boundary], a rational proper map of balls has no poles on the spheres, so we get a sphere map. Conversely a rational map taking a sphere to a sphere is a proper ball map via maximum principle. In [helal-2025-proper], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere. Putting all this together one gets the following equivalence.
Corollary 2.3.
Suppose $n\geq 2$. An $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper holomorphic map if and only if $f$ is the restriction of a rational map $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{n}$ with no poles on the spheres $rS^{2n-1}$ nor $S^{2n-1}$ such that $f(rS^{2n-1})\subset RS^{2N-1}$ and $f(S^{2n-1})\subset S^{2N-1}$.
If $n=N=1$ the rationality still holds (but not the bound for the degree) as we just showed above. However, if $N>n=1$, then nonrational maps exist.
Example 1.
Next, let us show that a nonrational proper map exists from $\mathbb{A}_{1,r}$ to $\mathbb{A}_{2,R}$ for some $R$. First, we construct a nonrational proper map from the disc $\mathbb{D}$ to the 2-ball $\mathbb{B}_{2}$. Such a map follows from the result of Dor [dor-1990-proper], but it is useful to consider an elementary construction in the one-dimensional case. Take an arbitrary nonrational function from the disc $f_{1}\colon\mathbb{D}\to\mathbb{D}$, suppose that it extends continuously to the closed disc $\overline{\mathbb{D}}$ but does not extend holomorphically past any point on $\partial\mathbb{D}$ and suppose that $\sup_{z\in\overline{\mathbb{D}}}\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert<1$. Find a real-valued harmonic function $u$ on $\mathbb{D}$ continuous on $\overline{\mathbb{D}}$ so that on $\partial\mathbb{D}$ we have $u(z)=\frac{1}{2}\log\mathopen{}\mathclose{{\left(1-\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}}}\right)$. Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$. Then write $f_{2}=e^{u+iv}$. Then the mapping $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$. Clearly $f$ is not spherically equivalent to a rational map if $f_{1}$ is not rational.
The map $f$ is at most points locally an embedding. We can compose with an automorphism on the source and target and assume that $f(0)=0$, $f^{\prime}(0)\not=0$, and that for some neighborhood $U\subset\mathbb{D}$ of 0 and a neighborhood $V\subset\mathbb{B}_{2}$ of 0, $f(U)\subset V$ is a properly embedded submanifold, $f|_{U}$ is one-to-one, and $f(\mathbb{D}\setminus U)\cap V=\varnothing$. Assume that $V$ is a ball. Consider a closed ball $B$ centered at the origin that is small enough so that $\overline{B}\subset V$ and so that $K=f^{-1}(B)$ is connected. The set $\mathbb{D}\setminus K$ is biholomorphic to some $\mathbb{A}_{1,r}$ for some $r$ via some $\varphi\colon\mathbb{A}_{1,r}\to\mathbb{D}\setminus K$. Then $f\circ\varphi$ is a proper map of $\mathbb{A}_{1,r}$ to $\mathbb{B}_{2}\setminus B=\mathbb{A}_{2,R}$ (where $R$ is the radius of $B$).
The set $\mathbb{D}\setminus K$ has real-analytic boundary and so $\varphi$ extends past the boundary. Therefore, $f\circ\varphi$ cannot extend past the boundary at any point on $\partial\mathbb{D}$ by the assumption on $f_{1}$ and so $f\circ\varphi$ is not rational.
Let us compare the effect of the automorphism groups for proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ versus $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form (9) below. For any affine subspace $L$ of dimension $m<n$ such that $L\cap\mathbb{B}_{n}\not$$\varnothing$, there exists a $\varphi\in\operatorname{Aut}(\mathbb{B}_{n})$ so that $\varphi(L)$ is given by $z_{m+1}=\cdots=z_{n}=0$.
The automorphism group of $\mathbb{A}_{n,r}$ is the unitary group. In particular, given an affine subspace $L$ of dimension $m<n$ that intersects both the outer ball $\mathbb{B}_{n}$ and the inner ball $r\mathbb{B}_{n}$, elementary geometry says that there is a $\varphi\in\operatorname{Aut}(\mathbb{A}_{n,r})=U(n)$ so that $\varphi(L)$ is given by $z_{m+1}=c$ and $z_{m+2}=\cdots=z_{n}=0$ where $0\leq c<r$.
The effect of this difference on classification up to equivalence with respect to the embedding dimension $N_{f}$ is the following. For a proper ball map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ where $N_{f}<N$, there exists an automorphism $\psi\in\operatorname{Aut}(\mathbb{B}_{N})$ so that $\psi\circ f=F\oplus 0$ where $F\colon\mathbb{B}_{n}\to\mathbb{B}_{N_{f}}$ is a proper map ($N_{f}=N_{F}$). The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space, it can only “rotate” it. The proof is an elementary computation.
Proposition 2.4.
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper map where $N_{f}<N$. Then there exists a unitary $U\in U(N)$ and a unique $c\in[0,R)$ such that
(8) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0$
where $F\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N_{f},\sqrt{\frac{R^{2}-c^{2}}{1-c^{2}}}}$ is a proper map and $N_{f}=N_{F}$.
In the proposition, the 0 in the direct sum has $N-N_{f}-1$ components (so it can be left out if $N=N_{f}+1$). The proposition leads to classification of first degree proper rational maps. We first start with the equidimensional case. In this case, unitaries (so unitarily equivalent to the identity) are the only maps.
Proposition 2.5.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ is a proper rational map of degree 1 and $n\geq 2$, then $r=R$ and $f$ is a unitary map.
Proof.
We have that $f$ is an LFT that must take the unit sphere to the unit sphere, hence it is well-known that it must have the form
(9) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z,$
where $U$ is a unitary and $a\in\mathbb{B}_{n}$. We need to show that $a=0$. By precomposing with a unitary, we can assume that $a=(a_{1},0,\ldots,0)$ and
(10) $\phi_{a}(z)=\mathopen{}\mathclose{{\left(\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}},\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{2}}{1-z_{1}\bar{a}_{1}},\ldots,\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{n}}{1-z_{1}\bar{a}_{1}}}}\right).$
We note that the complex line $z_{2}=\cdots=z_{n}=0$ gets taken by $\phi_{a}$ to this same complex line, and on this line $\phi_{a}$ is simply the corresponding LFT in one dimension $\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$.
It remains to show that if $z_{1}\mapsto\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$ takes a circle of radius $r<1$ centered at zero to a circle of radius $R<1$ centered at zero. In other words, we must have that for all $\theta$, $\mathopen{}\mathclose{{\left\lvert a_{1}-re^{i\theta}}}\right\rvert^{2}=R^{2}\mathopen{}\mathclose{{\left\lvert 1-re^{i\theta}\bar{a}_{1}}}\right\rvert^{2}$. By expanding and equating coefficients one obtains that $a_{1}=0$ and $r=R$. In other words, $a=0$ and $f$ is a unitary. ∎
As a quick corollary we can remove the hypothesis of rationality and degree. By Hartogs phenomenon and the maximum principle, we find that a holomorphic proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself. That this map is a rational map of degree 1 then follows by Alexander’s theorem.
Corollary 2.6.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ is a proper holomorphic map and $n\geq 2$, then $r=R$ and $f$ is a unitary map.
When $N>n$, the result is a little more complicated.
Proposition 2.7.
If $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree 1 and $N>n\geq 2$, then $R\geq r$ and $f$ is unitarily equivalent to the affine embedding
(11) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z\oplus\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}\oplus 0.$
In the proof we will use the notion of a rank of a real polynomial $r(z,\bar{z})$ in $\mathbb{C}^{n}$. By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
(12) $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
where $\epsilon_{j}\pm 1$ and $P_{j}$ are holomorphic polynomials.
Proof.
Write $f=\frac{p}{q}$. As $f$ takes the unit sphere to the unit sphere we have
(13) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert q(z)}}\right\rVert^{2}=\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$
where $\lambda$ has to be a constant as $p$ and $q$ are of degree 1. Therefore, the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ has to be $n$ as that is the number of positive eigenvalues of the matrix of coefficients $\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$. In other words, $N_{f}=n$ and we can apply Proposition˜2.4. The $F$ in the proposition is unitarily equivalent to the identity by Proposition˜2.5. The values of the factors follow by elementary computation. ∎
We derive a few of basic results about the $k_{f}$ integer. First we note that for any rational map, it is enough to check an open dense subset of hyperplanes, that is, the general hyperplane. Similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication. The space of affine hyperplanes in $\mathbb{C}^{n}$ has the topology of the Grassmannian (minus the hyperplane at infinity).
Proposition 2.8.
If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map, then $\dim\operatorname{aff}\bigl(f(H)\bigr)=k_{f}$ for an open dense set of affine hyperplanes $H$.
The proof is to simply note that the affine dimension of $f(H)$ must be lower semicontinuous on the space of hyperplanes $H$, since the set of hyperplanes where the affine dimension is at most some dimension is defined by polynomial equations.
Next, we recall the proof of the classic result that $k_{f}\leq N-1$ for rational maps of spheres, as the technique will be useful later.
Proposition 2.9.
If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
Proof.
Let $\langle\cdot,\cdot\rangle$ denote the standard sesquilinear inner product so that the sphere $S^{2n-1}$ is defined by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}=\langle z,z\rangle=1$. For any fixed $w\in\mathbb{C}^{n}$, define the Segre variety
(14) $\Sigma_{w}=\{z\in\mathbb{C}^{n}:\langle z,w\rangle=1\}.$
Similarly if $\omega\in\mathbb{C}^{N}$, we have the Segre variety of the target sphere $\Sigma_{\omega}\subset\mathbb{C}^{N}$. Note that the Segre variety for a sphere is an affine hyperplane. If $f=\frac{p}{q}$, then the fact that $f$ takes the sphere to the sphere means that $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}=g(z,\bar{z})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1}}\right)$, or $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}-1=\frac{g(z,\bar{z})}{\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1}}\right)$, where $g$ is a polynomial. By polarizing this identity, we find the equality
(15) $\mathopen{}\mathclose{{\left\langle f(z),f(w)}}\right\rangle-1=\frac{g(z,\bar{w})}{q(z)\overline{q(w)}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right).$
Writing $P=q^{-1}(0)$ for the pole set of $f$, we therefore have that $f(\Sigma_{w}\setminus P)\subset\Sigma_{f(w)}$ for all $w$ where $f(w)$ is defined. As any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$. Therefore, there is an open set of affine hyperplanes which $f$ takes into hyperplanes. Therefore $k_{f}\leq N-1$. ∎
The $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
Proposition 2.10.
For a map $f\colon U\subset\mathbb{C}^{n}\to\mathbb{C}^{N}$, $\phi$ an LFT on $\mathbb{C}^{n}$, and $\psi$ an LFT on $\mathbb{C}^{N}$, we have $k_{f}=k_{\psi\circ f\circ\phi}$. In other words, $k_{f}$ is both a spherical and a unitary invariant.
As part of Theorem˜1.2, we claim that $k_{H_{d}}=N-1$. Let us prove that now.
Proposition 2.11.
For the homogeneous map $H_{d}$, we have $k_{H_{d}}=N-1$ where $N=\binom{n+d-1}{d}$ is the target dimension of $H_{d}$.
Proof.
The map $H_{d}$ has as its components every degree $d$ monomial in $n$ variables, $H_{d}=(z_{1}^{d},\sqrt{d}\,z_{1}^{d-1}z_{2},\ldots)$. If we consider what $H_{d}$ does to the hyperplane $\{z_{1}=c\}$ for nonzero constants $c$, we find that except for the first component which becomes the constant $c^{d}$, all the other components are affine independent. In other words, the affine span of $H_{d}(\{z_{1}=c\})$ is of dimension N1N-1. ∎
Next we consider the $k_{f}$ integer for juxtapositions. We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [dangelo-2009-complexity]. The next proposition therefore shows that for rational ball (or annulus) maps, $k_{f}=N_{f}-1$ happens only on the non-extremal points of this convex set.
Proposition 2.12.
Suppose that $F_{t}$ is a juxtaposition of two rational proper maps of balls (or annuli) $F_{0}$ and $F_{1}$, that is, $F_{t}=\sqrt{1-t}\,F_{0}\oplus\sqrt{t}\,F_{1}$ for some $t\in(0,1)$, and further suppose that $\mathopen{}\mathclose{{\left\lVert F_{0}(z)}}\right\rVert^{2}\not\equiv\mathopen{}\mathclose{{\left\lVert F_{1}(z)}}\right\rVert^{2}$. Then $k_{f}\leq N_{f}-2$.
The condition on the norms is to ensure that we are taking the convex combination of a point with itself in the space of squared norms. The condition is equivalent to requiring there is no unitary $U$ such that $F_{1}=UF_{0}$ (possibly adding zero components to make the target dimensions equal).
Proof.
We note that if $k_{F_{0}}\leq N_{F_{0}}-2$ or $k_{F_{1}}\leq N_{F_{1}}-2$, then we are done trivially. Therefore suppose that $k_{F_{0}}=N_{F_{0}}-1$ and $k_{F_{1}}=N_{F_{1}}-1$. Also note from the proof of Proposition˜2.9 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies. In particular, we find these vectors in an $N_{f}$-dimensional hyperplane which we can assume is the entire target space. Apart from these facts that we can simply treat these as rational maps. The integers $k_{f}$ and $N_{f}$ and the fact about the normal vectors are all invariant under any LFT and so we only need to prove the result for $F_{0}\oplus F_{1}$. We can also assume that $F_{0}$ has $N_{F_{0}}$ components and $F_{1}$ has $N_{F_{1}}$ components. We can further homogenize both maps up to the same degree and instead of affine hyperplanes we deal with linear subspaces. The dimensions increase by 1, but that is not relevant to us as we are interested in their difference. In other words, what we have is two homogeneous polynomial maps of the same degree $F_{0}$ and $F_{1}$, we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent. We further have that for any codimension 1 subspace $H$, the span of $F_{0}(H)$ is of codimension 1 (in the codomain of $F_{0}$) and same for $F_{1}(H)$. It is possible that components of $F_{0}$ and $F_{1}$ are not linearly independent when put together. After composing with a linear we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components. Write the target space as $\mathbb{C}^{j}\times\mathbb{C}^{k}\times\mathbb{C}^{\ell}$. The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ ($v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, $0\in\mathbb{C}^{\ell}$) so that $\xi\cdot v=0$ for all $\xi\in P(H)$. There is also a vector $w=(w_{1},0,w_{3})$ so that $\xi\cdot w=0$ for all $\xi\in P(H)$. We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, we found that $P(H)$ lives in a codimension 2 space and we are done. Thus suppose that there exists a $v=w=(v_{1},0,0)$ for every subspace $H$. But our assumption about the normal vectors of the hyperplanes would give a contradiction. Thus $v$ and $w$ are always linearly independent and we find that $N_{P}-k_{p}\geq 2$. ∎
We have seen that the homogeneous map $H_{d}$ is an annulus map. We also have that juxtapositions of homogeneous maps and their affine embeddings are annulus maps. Here is an example that is not constructed in this way from homogeneous maps.
Example 2.
With the construction introduced in [helal-2025-proper], one finds the following map
(16) $f(z_{1},z_{2})=\mathopen{}\mathclose{{\left(\tfrac{2}{\sqrt{5}}z_{1}^{3},\tfrac{2\sqrt{2}}{\sqrt{5}}z_{1}^{2}z_{2},\tfrac{2}{\sqrt{5}}z_{1}z_{2}^{2},z_{1}z_{2},z_{2}^{2},\tfrac{1}{\sqrt{5}}z_{1}}}\right).$
This map gives a proper map of $\mathbb{A}_{2,\frac{1}{2}}$ to $\mathbb{A}_{6,\frac{1}{4}}$. It takes the unit sphere to the unit sphere and the half sphere to quarter sphere. It takes no other sphere centered at the origin to a sphere centered at the origin. It is not unitarily equivalent to a juxtaposition of several homogeneous maps, nor is it an affine embedding of such a map. If it were, then it would take all spheres centered at the origin to spheres centered at the origin. We have that $N_{f}=6$ and one can compute that the $k_{f}$ number is $3$, so $k_{f}<N_{f}-1$. To see why it is $3$, note that if we restrict to a 1-dimensional affine space $H$, the map becomes a one variable cubic polynomial and so the affine span can be at most $3$. That it is exactly $3$ can be seen by restricting to a hyperplane making $z_{2}$ constant.
We remark that while Theorem˜1.2 implies that for annuli maps, $k_{f}=N-1$ means $f$ is homogeneous, there must be extremal points with $k_{f}<N-1$ in the convex set describing the rational proper annulus maps as the example above shows that there exist maps that are not constructed out of homogeneous maps via juxtapositions.
The construction from [helal-2025-proper] can be used to obtain cubic examples in higher dimensions and also quadratic rational maps with the same properties (not taking more than two zero-centric spheres to zero-centeric spheres, so not constructed out of homogeneous maps).
3. First gap
Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [dangelo-2012-pfisters]*Theorem 1.1. We only need the quadratic version of this result, but we state the more general result. Recall that the rank of a real polynomial $r(z,\bar{z})$ is the rank of its matrix of coefficients, or alternatively the smallest $k$ such that $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2}$.
Theorem 3.1.
Let $n\geq 1$, $s_{1},s_{2}$ nonzero real numbers of the same sign, and $Q(z,\bar{z})$ be a nonzero real polynomial on $\mathbb{C}^{n}$. Then
(17) $\operatorname{rank}Q(z,\bar{z})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})\geq\operatorname{rank}(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{1})(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}+s_{2})=\binom{n+2}{2}$.
The proof follows by roughly the same argument as in [dangelo-2012-pfisters] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors. What must be newly proved here is a “diagonal version” of the theorem. We only have this proof for two factors, which is sufficient for our purposes. We conjecture that the result holds for any number of factors. Note that the result is not true as stated if the numbers $s_{j}$ have opposite signs. For example, $(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}+1)(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}-1)=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{4}-1$ is of rank 2 and not $\binom{n+2}{2}=\binom{3}{2}=3$.
Proof.
The main new claim that must be proved is the “diagonal version” of this theorem. That is, for a real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
(18) $\bigl(Q(x)(x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)\\ \geq\#\bigl((x_{1}+\cdots+x_{n}+s_{1})(x_{1}+\cdots+x_{n}+s_{2})\bigr)=\binom{n+2}{2}$,
where #\# means the number of nonzero coefficients in the given polynomials when they are expanded. We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{d}$ and applying the claim for positive $s_{1},s_{2}$.
First suppose that $n=1$. For this case, assume $s_{1}$ and $s_{2}$ are both negative. The polynomial
(19) $R(x)=Q(x)(x+s_{1})(x+s_{2})$
has at least $2$ positive real roots. By Descartes’ rule of signs, there must be at least $2$ sign changes among the coefficients of the polynomial and therefore there must be at least $3$ nonzero coefficients, which is precisely the claimed result for $n=1$.
Next, let us suppose $n=2$. Again let $R(x)=Q(x)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$. Write $R(x)=x_{1}^{j}x_{2}^{k}\widetilde{R}(x)$ (and similarly $Q(x)=x_{1}^{j}x_{2}^{k}\widetilde{Q}(x)$) where $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ are both not identically zero. Applying, the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively). One of these terms can be the constant, and therefore, we have $1+2n$ terms in $\widetilde{R}(x)$ which are all pure powers. Next we will show that there is a term in $\widetilde{R}(x)$ that involves both $x_{1}$ and $x_{2}$. Note that the vector field $L=\frac{\partial}{\partial x_{1}}-\frac{\partial}{\partial x_{2}}$ is tangent to each $(x_{1}+x_{2}+s_{j})$. Therefore, if we apply $L$ $\nu$ number of times, we get
(20) $L^{\nu}\widetilde{R}(x)=\bigl(L^{\nu}\widetilde{Q}(x)\bigr)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2}).$
As $\widetilde{Q}$ is a polynomial, $L^{\nu}\widetilde{Q}$ must be identically zero for some $\nu$, so take $\nu$ such that $L^{\nu-1}Q$ is not zero but $L^{\nu}Q\equiv 0$. Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation and so it is a function $x_{1}+x_{2}$, that is $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$ $S$ is not identically zero. But then
(21) $L^{\nu-1}\widetilde{R}(x)=S(x_{1}+x_{2})(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$
is a polynomial in $(x_{1}+x_{2})$ of degree at least 2 and hence it must have a mixed monomial. Since $L^{\nu-1}\widetilde{R}(x)$ has a mixed monomial, then so does $\widetilde{R}(x)$ as $L$ cannot manufacture mixed monomials out of pure ones. Therefore, $R(x)$ has at least $1+2\cdot 2+1=\binom{2+2}{2}=6$ terms.
Moreover, we guarantee that there are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$, and at least 3 terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m\geq k$. Note that also there are no terms at all with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$. Similarly, there are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$, and at least 3 terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell>k$ and $m\geq j$. And no terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell<k$.
We now attack the general problem for $n>2$. Pick any pair of variables, without loss of generality it is $x_{1}$ and $x_{2}$. Write $x=(x_{1},x_{2},x^{\prime})$ and fix one of the standard monomial orders on the remaining variables $x^{\prime}$, such that multiplying a monomial in $x^{\prime}$ by any one of the variables $x_{3},\ldots,x_{n}$ moves the monomial further in the ordering. Using multinomial notation, write
(22) $R(x)=\sum_{\alpha}r_{\alpha}(x_{1},x_{2})(x^{\prime})^{\alpha}\quad\text{and}\quad Q(x)=\sum_{\alpha}q_{\alpha}(x_{1},x_{2})(x^{\prime})^{\alpha}.$
we find the smallest monomial $(x^{\prime})^{\mu}$ in the given monomial order for which the $q_{\alpha}$ is not identically zero. We then have that
(23) $r_{\mu}(x_{1},x_{2})=q_{\mu}(x_{1},x_{2})(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2}).$
All the monomials of $r_{\mu}(x_{1},x_{2})(x^{\prime})^{\mu}$ must appear in $R(x)$ by minimality of $\mu$. We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these $a_{\ell m}$ is as in the “Moreover” remark above. We use this this remark for every pair of variables to count monomials. For every pair there is one monomial that we might be also counting for other pairs, so we count it only once. Then there are two monomials corresponding to each variable that may appear in the count for every pair that includes this variable. Finally, for every pair there is at least one monomial that does not appear in the count for every other pair. In other words, We have at least
(24) $1+2n+\binom{n}{2}=\binom{n+2}{n}$
distinct monomials in $R(x)$. The claim is proved.
Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [dangelo-2012-pfisters]. That is, there exists some ordering of monomials so that if we write out the matrix of coefficients of the expressions involved, multiplication by $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ moves along diagonals. The rank of the matrix of coefficients is the rank of the function. The rank of a matrix can be estimated from below by considering the number of nonzero terms in the superdiagonal that is furthest up and to the right that has nonzero terms. Up to multiplication with a monomial, the terms on this diagonal are polynomials in $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$, and so by replacing $x_{j}=\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ we obtain the situation of the claim and the proof follows. See [dangelo-2012-pfisters] for more details. ∎
We can now prove Theorem˜1.1, that is, if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ ($n\geq 2$) is a proper holomorphic map and $n<N<\binom{n+1}{2}$, then $f$ is unitarily equivalent to an affine embedding (3) and $r\leq R<1$.
Proof of Theorem˜1.1.
By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$. Lemma 3.11 from [helal-2025-proper] gives us
(25) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\Bigr)\cdot\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q_{2}(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
for a real polynomial $Q_{2}(z,\bar{z})$.
If $Q_{2}\equiv 0$, then by the argument in the proof of Theorem 3.12 from [helal-2025-proper], $f$ is a polynomial, that is, $q\equiv 1$. Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$). Therefore, $\deg f=1$. Proposition˜2.7 implies that $f$ is an affine embedding and that $R\geq r$.
Suppose $Q_{2}\not\equiv 0$, we will get a contradiction. By applying Proposition˜2.4, we reduce to the case when $N=N_{f}$. The dimension $N$ is then the number of components of $f$, which is equal to the number of components of $p$, which we can assume are linearly independent as $N_{f}=N$. Therefore, $N$ is equal to the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$. Theorem˜3.1 gives us
[…]
In the inequalities we have used the facts that $\operatorname{rank}(A+B)\geq-(\operatorname{rank}A)+\operatorname{rank}B$ and that $\operatorname{rank}\bigl(\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}A(z,\bar{z})\bigr)=\operatorname{rank}A(z,\bar{z})$, both of which follow immediately from the definition of rank. By hypothesis, we find a contradiction as we assumed $N_{f}<\binom{n+1}{2}$.
We therefore obtained $N_{f}=n$ and that the degree is $1$. The conclusion follows from Proposition˜2.5 or Proposition˜2.7 depending on whether $N>n$ or $N=n$. The interval is sharp; when $N=\binom{n+1}{2}$, then besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding. ∎
We should remark that for specific $\mathbb{A}_{n,r}$ and $\mathbb{A}_{\binom{n+1}{2},R}$ we either get the affine embeddings if $R\geq r$, or we get the homogeneous map $H_{2}$ if $R=r^{2}<r$. For general $n$, it is not clear if other maps exist when $N=\binom{n+1}{2}$, for perhaps different $R$. For $n=2$, the homogeneous $H_{2}$ and the affine embeddings are the only proper maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{\binom{2+1}{2},R}=\mathbb{A}_{3,R}$. See Theorem˜1.4.
4. Homogeneous maps via the generic hyperplane rank
To prove the theorem on classification of homogeneous maps, we will prove a more general result about rational maps taking two spheres to spheres. To state the result, it may be easiest to work in the projective space $\mathbb{P}^{n}$. The automorphism group $\operatorname{Aut}(\mathbb{P}^{n})$ of $\mathbb{P}^{n}$ is the set of linear fractional transformations or LFTs. Consider some embedding $\mathbb{C}^{n}\subset\mathbb{P}^{n}$ and then we can consider the sphere $S^{2n-1}$ as a subset of $\mathbb{P}^{n}$. When we refer to affine coordinates, we refer to this embedding and the coordinates $(z_{1},\ldots,z_{n})$. We say $\mathcal{S}$ is an LFT sphere if there exists an LFT $\varphi\in\operatorname{Aut}(\mathbb{P}^{n})$ so that $\mathcal{S}=\varphi(S^{2n-1})$.
Our first result is about rational maps taking two LFT spheres to LFT spheres. The definition of $k_{f}$ for a rational map $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in the projective setting is exactly analogous to the affine setting; it is the smallest dimension $k$ of a $k$-plane in $\mathbb{P}^{N}$ that contains $f(H)$ for a general hyperplane $H\subset\mathbb{P}^{n}$. In the proof we will work in homogeneous coordinates on the projective space, and so a general hyperplane in $H\subset\mathbb{P}^{n}$ is an $n$-dimensional linear subspace in $\mathbb{C}^{n+1}$.
We use $z=(z_{0},\ldots,z_{n})$ for the homogeneous coordinates on $\mathbb{P}^{n}$. To write a rational function $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ in homogeneous coordinates we start with $f=\frac{(p_{1},\ldots,p_{N})}{q}$ and order it as $(q,p_{1},\ldots,p_{N})$ and homogenize, that is, we make the denominator the 0th component.
The sphere $S^{2n-1}\subset\mathbb{C}^{n}\subset\mathbb{P}^{n}$ in our chosen affine embedding when put into homogeneous coordinates is given by the equation
(27) $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0,\quad\text{where}\quad J_{n}=\operatorname{diag}(-1,1,\ldots,1)=\begin{bmatrix}-1&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
[…]
(28) $\mathopen{}\mathclose{{\left\langle T^{*}J_{n}Tz,z}}\right\rangle=0,$
for an invertible $T$, that is, it is $\mathopen{}\mathclose{{\left\langle Az,z}}\right\rangle=0$ for some hermitian $(n+1)\times(n+1)$ matrix $A$ with $n$ positive and $1$ negative eigenvalues. If $\varphi$ is the LFT denoted by the matrix $T$, then $\mathopen{}\mathclose{{\left\langle T^{*}J_{n}Tz,z}}\right\rangle=0$ represents the sphere $\varphi^{-1}(S^{2n-1})$. We can also do the same in $\mathbb{C}^{N+1}$ on the target, using instead the matrix $J_{N}$.
Theorem 4.1.
Suppose $f\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$ is a rational map that takes two distinct LFT spheres to two LFT spheres and $k_{f}=N-1$. Then $f\circ\alpha=\beta\circ f$ for some LFTs $\alpha\in\operatorname{Aut}(\mathbb{P}^{n})$ and $\beta\in\operatorname{Aut}(\mathbb{P}^{N})$ neither of which is the identity.
More precisely, using the coordinates above, if one of the spheres on the source (resp. the target) are given by $J_{n}$ (resp. $J_{N}$) and the other by a hermitian matrix $A$ (resp. $B$), then $\alpha$ is represented by the matrix $J_{n}A$ and $\beta$ is represented by the matrix $J_{n}B$.
The main idea behind the proof is that $k_{f}=N-1$ means that the map $f$ induces a map of the corresponding Grassmannians. To figure out the map one uses the Segre variety of the LFT spheres. But now we have two LFT spheres that go to LFT spheres and the second LFT sphere pair means that $f$ induces a different mapping of Grassmannians. Comparing the two maps gives us the corresponding $\alpha$ and $\beta$.
Proof.
As we said we will work in homogeneous coordinates, that is, in $\mathbb{C}^{n+1}$ and $\mathbb{C}^{N+1}$. Starting with a rational map $f$, we clear denominators and homogenize to represent $f$ by a homogeneous polynomial map $F=(F_{0},F_{1},\ldots,F_{N})\colon\mathbb{C}^{n+1}\to\mathbb{C}^{N+1}$ where $F_{0}$ is the denominator. After applying an LFT we can assume that one of the matrices on both the source and the target is the standard sphere $S^{2n-1}$ and $S^{2N-1}$, that is, it is given by $\mathopen{}\mathclose{{\left\langle J_{n}z,z}}\right\rangle=0$. For any hermitian matrix $M$ and a point $w$, we write the Segre variety of the homogenized LFT sphere given by $\mathopen{}\mathclose{{\left\langle Mz,z}}\right\rangle=0$ as
(29) $\Sigma_{w}^{M}=\{z:\mathopen{}\mathclose{{\left\langle Mz,w}}\right\rangle=0\}.$
In affine coordinates, the fact that $f=\frac{p}{q}$ takes the standard sphere to the standard sphere means $\mathopen{}\mathclose{{\left\langle p(z),p(w)}}\right\rangle-q(z)\overline{q(w)}=g(z,\bar{w})\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle z,w}}\right\rangle-1}}\right)$ for a polynomial $g$. In homogeneous coordinates, this translates to
(30) $\mathopen{}\mathclose{{\left\langle J_{N}F(z),F(w)}}\right\rangle=G(z,\bar{w})\mathopen{}\mathclose{{\left\langle J_{n}z,w}}\right\rangle.$
This equation implies that
(31) $F(\Sigma_{w}^{J_{n}})\subset\Sigma_{F(w)}^{J_{N}}.$
We are assuming that we have a homogenized LFT sphere in $\mathbb{C}^{n+1}$ represented by the $(n+1)\times(n+1)$ hermitian matrix $A$ that $F$ takes to a homogenized LFT sphere in $\mathbb{C}^{N+1}$ represented by the $(N+1)\times(N+1)$ hermitian matrix $B$. Therefore, we also have
(32) $F(\Sigma_{w}^{A})\subset\Sigma_{F(w)}^{B}.$
Let $S=(AJ_{n}^{-1})^{*}=J_{n}A$. As
(33) $\mathopen{}\mathclose{{\left\langle Az,w}}\right\rangle=\mathopen{}\mathclose{{\left\langle AJ_{n}^{-1}J_{n}z,w}}\right\rangle=\mathopen{}\mathclose{{\left\langle J_{n}z,(AJ_{n}^{-1})^{*}w}}\right\rangle,$
we find that $\Sigma_{w}^{A}=\Sigma_{Sw}^{J_{n}}$, and so
(34) $F(\Sigma_{w}^{A})\subset\Sigma_{F(w)}^{B}\quad\text{and}\quad F(\Sigma_{w}^{A})=F(\Sigma_{Sw}^{J_{n}})\subset\Sigma_{F(Sw)}^{J_{N}}.$
The condition that $k_{f}=N-1$ means in homogeneous coordinates that given a linear $n$-dimensional subspace $V\subset\mathbb{C}^{n+1}$, the linear span of $F(V)$ in $\mathbb{C}^{N+1}$ is $N$-dimensional. The Segre variety $\Sigma_{w}^{A}$ is an $n$-dimensional subspace, and $\Sigma_{F(w)}^{B}$ and $\Sigma_{F(Sw)}^{J_{N}}$ are $N$-dimensional subspaces. In other words, we find that
(35) $\Sigma_{F(Sw)}^{J_{N}}=\Sigma_{F(w)}^{B}.$
That is, using coordinates $(Z_{0},\ldots,Z_{N})$ in $\mathbb{C}^{N+1}$, we have that $\mathopen{}\mathclose{{\left\langle J_{N}Z,F(Sw)}}\right\rangle=0$ and $\mathopen{}\mathclose{{\left\langle BZ,F(w)}}\right\rangle=0$ denote the same subspace. As this is true for every $w$, we have that $\mathopen{}\mathclose{{\left\langle J_{N}Z,F(Sw)}}\right\rangle=\lambda\mathopen{}\mathclose{{\left\langle BZ,F(w)}}\right\rangle$, where $\lambda$ must be a constant as the two functions are both degree-$1$ homogeneous in $Z$ and degree-$d$ homogeneous in $w$ (assuming $F$ is of degree $d$). We can rewrite this equation as $J_{N}^{*}F(Sw)=\lambda B^{*}F(w)$, or in other words,
(36) $F(Sw)=\lambda J_{N}^{-*}B^{*}F(w).$
Note that $J_{N}^{-*}=J_{N}$ and $B=B^{*}$, so let $T=J_{N}^{-*}B^{*}=J_{N}B$ and then we have that for all $w$, we have
(37) $F(Sw)=\lambda TF(w).$
If we take $\alpha$ to be the LFT given by the matrix $S$ and $\beta$ be the LFT given by the matrix $T$, then we obtain in affine coordinates that
(38) $f\circ\alpha=\beta\circ f,$
as required. The formulas for $\alpha$ and $\beta$ follow. ∎
In particular, $f$ now also takes the LFT sphere $\alpha^{-1}(S^{2n-1})$ to the LFT sphere $\beta^{-1}(S^{2N-1})$. It turns out that these spheres need not be the ones we started with. By iterating, we find that for any $k\in\mathbb{Z}$ we have that
(39) $f\circ\alpha^{k}=\beta^{k}\circ f.$
So we also get that $f$ takes $\alpha^{-k}(S^{2n-1})$ to $\beta^{-k}(S^{2N-1})$ for any $k$. In our particular case, when all spheres involved are centered at zero, $f$ takes infinitely many spheres centered at zero of various radii to spheres centered at zero. In particular, in the first step, whenever it takes the standard sphere to the standard sphere and the $r$-sphere to the $R$-sphere, then we find that it also takes the $r^{2}$-sphere to the $R^{2}$-sphere and so on.
Lemma 4.2.
Suppose $n\geq 2$ and $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map such that $f(S^{2n-1})\subset S^{2N-1}$ and for some $r,R<1$ we have that $f(rS^{2n-1})\subset RS^{2N-1}$. Suppose also that $k_{f}=N-1$. Then there are two sequences of distinct numbers $\{r_{k}\}$ and $\{R_{k}\}$ converging to zero so that $f(r_{k}S^{2n-1})\subset R_{K}S^{2N-1}$ for every $k$.
Proof.
We set things up as in the proof of Theorem˜4.1, and so the matrices $A$ and $B$ are given as
(40) $A=\begin{bmatrix}-r^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}\quad\text{and}\quad B=\begin{bmatrix}-R^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
Then
(41) $S=J_{n}A=\begin{bmatrix}r^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}\quad\text{and}\quad T=J_{N}B=\begin{bmatrix}R^{2}&0&\cdots&0\\ 0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{bmatrix}$.
As we said what we have is that $f$ takes $\alpha^{-1}(S^{2n-1})$ to $\beta^{-1}(S^{2N-1})$. In affine coordinates, $\alpha$ is simply a scaling by $\frac{1}{r^{2}}$ and $\beta$ is a scaling by $\frac{1}{R^{2}}$, and so $\alpha^{-1}(S^{2n-1})=r^{2}S^{2n-1}$ and $\beta^{-1}(S^{2N-1})=R^{2}S^{2N-1}$. By iterating we get the desired sequence. ∎
We are now ready to prove Theorem˜1.2 using the results of our previous work in [helal-2025-proper].
Proof of Theorem˜1.2.
One direction of the theorem is clear and follows from the discussion so far. What is left to prove is that if $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a holomorphic proper map and $k_{f}=N-1$, then $f=UH_{d}$ for some unitary $U$ and an integer $d$. By Theorem˜2.2, $f$ is rational and takes spheres to spheres as in Lemma˜4.2. Therefore, $f$ takes infinitely many spheres centered at zero to spheres centered at zero. In [helal-2025-proper], we have proved that such a rational map must, up to a postcomposition with a unitary, be a direct sum of scaled maps $H_{d}$ for various $d$ or zeros. So for some unitary $U$, we can assume that
(42) $f=U(c_{1}H_{d_{1}}\oplus\cdots c_{\ell}H_{d_{\ell}}\oplus 0),$
where we can assume that $d_{1},\ldots,d_{\ell}$ are distinct. Since we are assuming that $k_{f}=N-1$, we must have that $N_{f}=N$ and hence there are no zero components. Then, via the argument in the proof of Proposition˜2.11, we can see that $k_{f}=N$-$\ell$. Hence $\ell=1$, and $f=UH_{d}$ as required. ∎
Let us now prove Theorem˜1.4 as an application of Theorem˜1.2. We already have that proper holomorphic maps from $\mathbb{A}_{2,r}$ to $\mathbb{A}_{3,R}$ are rational. If the degree of the map is 1, then via Proposition˜2.7, it is unitarily equivalent to the affine embedding
(43) $(z_{1},z_{2})\mapsto\mathopen{}\mathclose{{\left(\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{1},\sqrt{\frac{1-R^{2}}{1-r^{2}}}\,z_{2},\sqrt{\frac{R^{2}-r^{2}}{1-r^{2}}}}}\right),$
where $R\geq r$. In this case, clearly $N_{f}=2$ and $k_{f}=1$.
The theorem will therefore follow if we can show that any map of degree greater than 1 is unitarily equivalent to $H_{2}$. This result follows by Faran’s [faran-1982-maps] classification of proper maps of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The following lemma finishes the proof of Theorem˜1.4.
Lemma 4.3.
Suppose that $f\colon\mathbb{A}_{2,r}\to\mathbb{A}_{3,R}$ is a rational proper map of degree greater than 1. Then $f$ is unitarily equivalent to the homogeneous map $H_{2}$.
Proof.
The map $f$ extends to a proper map of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. Faran’s result implies that a proper rational map of $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$ is spherically equivalent to one of four maps: the linear embedding $(z_{1},z_{2},0)$, the Whitney map $(z_{1},z_{1}z_{2},z_{2}^{2})$, the homogeneous $(z_{1}^{2},\sqrt{2}\,z_{1}z_{2},z_{2}^{2})$, and the map $(z_{1}^{3},\sqrt{3}\,z_{1}z_{2},z_{2}^{3})$. The issue is that the automorphisms used in Faran’s classification are automorphisms of the ball and these do not preserve the inner sphere boundary of the annulus. Spherical equivalence does, however, preserve the degree of the map and the integer $k_{f}$. For the three maps of degree greater than 1, we find $k_{f}=2=N-1$. To compute $k_{f}$, it is sufficient to consider the hyperplanes $z_{2}=\text{constant}$. In other words, if we have an annulus map as in the statement and of degree greater than 1, then $k_{f}=2=N-1$. Hence Theorem˜1.2 says that $f$ is unitarily equivalent to $H_{2}$. ∎
5. Other remarks on proper maps of annuli
Let us put together some further minor remarks. Firstly, there are two other domains in $\mathbb{C}^{n}$ whose automorphism group is $U(n)$. There is the punctured ball $\mathbb{B}_{n}\setminus\{0\}$ and the complement of the closed unit ball $\mathbb{C}^{n}\setminus\overline{\mathbb{B}_{n}}$. For maps from a punctured ball to a punctured ball, the maps must extend through the origin by the Riemann extension theorem if $n=1$ or via the Hartogs phenomenon if $n>1$. We are then simply dealing with proper maps of balls that take the origin to the origin. Since the automorphism group of the ball is transitive, every spherical equivalence class of proper maps of balls contains a map taking 0 to 0. The case of the complement of the closed unit ball was studied in [helal-2025-proper]. In that case, for $n\geq 2$, the proper maps are polynomial sphere maps whose norm goes to infinity at infinity. In either case, the classification is different in these degenerate cases, and in both cases there are more maps than for annuli.
Another remark we wish to make is that we have already seen that in the first interval we have the restriction $r\leq R<1$ for a map to exist. A question may arise about which $R$ admit a proper map in general. It is relatively simple to see that as long as we do not place any restriction on $N$ we may get any $r$ and $R$ whatsoever. That is, given $n$, $r$, and $R$, pick $d$ so that $r^{d}\leq R$. Then as $H_{d}$ takes the $r$-sphere to the $r^{d}$-sphere we have that the map
(44) $z\mapsto\sqrt{\frac{1-R^{2}}{1-r^{2d}}}\,H_{d}(z)\oplus\sqrt{\frac{R^{2}-r^{2d}}{1-r^{2d}}}$
is a proper map of $\mathbb{A}_{n,r}$ to $\mathbb{A}_{N,R}$ where $N=\binom{n+d-1}{d}+1$.
The previous discussion suggests that the existence of a rational proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ of degree $d$ may imply a relationship between $r$, $R$, and $d$. For homogeneous maps as above we have $r^{d}\leq R$, with $H_{d}$ itself achieving the equality in this bound. Juxtapositions (that is, weighted direct sums) of homogeneous maps, and of all the other maps the authors have constructed still satisfy this bound, and so a reasonable question seems if $r^{d}\leq R$ (equivalently as a bound on degree, $d\geq\log_{r}R$) holds in general. If one also requires that $N$ is the embedding dimension ($N=N_{f}$), then more can be said. In this case, if $d=1$, then $R=r$. For higher degree, one could use juxtaposition to get a range. For example, if we juxtapose $H_{1}$ and $H_{d}$, then via the calculation in the introduction, we have that for $t\in(0,1)$, the map $\sqrt{1-t}\,H_{1}\oplus\sqrt{t}\,H_{d}$ has $R=\sqrt{(1-t)r^{2}+tr^{2d}}$, and so any $R$ in the range between $r^{d}$ and $r$ is achieved by these maps. So the question could be modified for the case when $N=N_{f}$ and degree $d>1$, to ask if $R$ equal to or bigger than $r$ is possible.
Finally, we remark that given our results and our examples, it is plausible to conjecture a degree bound in terms of the dimensions. Some bound must exist by the theorem of Forstnerič, as annulus maps extend to proper rational maps of balls. However, in the annulus case, it appears that the homogeneous map achieves the highest degree for a given set of dimensions. That is, we conjecture that as long as $n\geq 2$, then $N\geq\binom{n+d-1}{d}$; note that this is a bound for $d$, but it does not have a nice formula. All the other maps we know how to construct are either juxtapositions of homogeneous maps or have lower degree than the homogeneous for a given target dimension.